Oil leaks from a tank. At hour t = 0 there are 220
Answer #1

100 At t=0 oil = 220 gallons Oil leaked in 1hr = 220 x3 = 6.6 I oil leaked in 6 hor= 6x6,6 = 39.6 Oil left affer 6 hor= 220-Olt) = ao ekt Aut to=0 90 = 220 At t=1 a(t) = 220-6.6 = 213.4 213,4 = 220 ekx! 213.4 ok 220=ek 0.97= ek ln 0.97= lnek -0.030